# Deployment and validation of predictive 6-dimensional beam diagnostics through generative reconstruction with standard accelerator elements

### v2

@@ Abstract
Particle accelerator-based user facilities are indispensable tools that power scientific discovery in the biological, chemical and material sciences [1, 2]. Understanding the precise distribution of particles in an accelerated beam is therefore critical to improving and maintaining optimal performance of these facilities. For example, X-ray Free Electron Laser (XFEL) facilities are strongly dependent on the distribution of electrons in 6-dimensional position-momentum space to generate high-brightness, ultra-short X-ray pulses [3, 4, 5, 6, 7, 8, 9, 10, 11]. The precise structure of the beam distribution during the acceleration and compression stages of the accelerator impacts overall beam quality due to collective effects such as space charge and coherent synchrotron radiation (CSR) [12, 13, 14, 15, 16, 17], that in-turn, negatively impacts FEL performance. As a result, methods are needed to efficiently and precisely characterize 6-dimensional phase space distribution at multiple locations along the beamline.
@@
While methods for measuring lower dimensional phase space distribution structure of particle beams are well understood, either through sub-sampling portions of the beam distribution [18, 19] or via tomographic manipulations [20, 21, 22, 23, 24], their applicability becomes limited on 5- [25, 26] or 6-dimensional [27] phase space distributions due to largely increased number of samples required by the nature of the tomography. Previous attempts at characterizing the high-dimensional phase space distribution of beams require a substantial amount of dedicated experimental beam time to conduct ($>$18 hrs for 6-dimensional characterization [27]), making these methods impractical for use at user facilities. Additionally, these methods also require the use of specialized diagnostic elements to characterize the beam distribution, such as movable slits or transverse deflecting cavities (TCAVs) [28, 29, 30, 31]. These additional requirements can add complexity and operational costs to diagnose the beam at multiple locations and reduce high-value space in the accelerator beamline.
Recently, machine learning methods have been used to create so-called “virtual diagnostics” that are able to predict the phase space distributions from beamline parameters and/or experimental measurements [32, 33, 34, 35, 36, 37, 38, 39]. These predictive models are trained on simulated data, where, for a given set of accelerator parameters, the true beam distribution at the location of the virtual diagnostic is known from beam physics principals. However, these methods rely on training data generated from accurate physics models of entire accelerator beamlines, which can become prohibitively difficult to effectively calibrate to match experimental accelerator conditions at the level of precision needed to resolve precise beam distributions.
Figure 1: a. Schematics of the GPSR method [40]. b: Cartoon showing an evolution of beam energy chirp by the accelerating field. c. Schematic view of the first bunch compressor section at the PAL-XFEL, where the longitudinal phase space is measured at YAG2 screen using S-band TCAV and spectrometer [e. measured longitudinal phase space for BC and XLIN OFF]. Nominal beam charge and measured kinetic energy are 250 pC and 260.5 MeV, respectively. The lattice used for the reconstruction training process is in the dashed orange box. d. Beam images at the YAG1 screen with various BC, XLIN, and Q1 conditions.
@@
More recently, a new machine learning technique, known as generative phase space reconstruction (GPSR), has been developed to address the limitations of high-dimensional characterization and predictive methods [40]. A schematic of the GPSR technique is as follows [see Fig. 1a]: the initial 6-dimensional phase space distribution is parameterized by a transformer-based, generative machine learning model that is able to flexibly generate a wide variety of distribution structures. The initially generated beam is then propagated through a simulated version of the beamline using a differentiable particle tracking software [41]. To reconstruct the beam distribution, the parameters of the generative model are iteratively updated via gradient descent optimization [42] to minimize the difference between experimental measurements of the $(x,y)$ beam distribution on intercepting screens and corresponding simulated predictions. Minimizing the difference ultimately yields the reconstructed beam phase space distribution that reproduces experimental measurements.
@@
Previous experimental demonstrations of the GPSR method [40, 43, 44] have shown a dramatic reduction in the number of experimental measurements needed to reconstruct 4 to 6-dimensional phase space distributions, leading to the corresponding reduction in the amount of beam time required to perform those measurements. Nonetheless, its broader applications at the accelerator facilities remains limited, as it still relies on specialized diagnostic elements such as TCAVs to rotate and probe the beam distribution in 6-dimensional phase space. In addition, prior studies have primarily focused on validating reconstructions at a specific diagnostic point, leaving predictive accuracy and uncertainty quantification at additional downstream locations unexplored.
@@
In this work, we demonstrate the first implementation and validation of the GPSR methodology for complete 6-dimensional phase space reconstruction using standard beamline elements commonly available at accelerator facilities [2, 45, 46], including accelerating cavities and spectrometer magnets. Through both simulation and experiment at the user facility, Pohang Accelerator Laboratory X-ray Free Electron Laser (PAL-XFEL) [47, 10, 11], we show that a combination of quadrupole scans, dispersive measurements, and the accelerating cavity phase scans provides sufficient information to reconstruct the detailed 6-dimensional phase space distribution of the beam including complex and nonlinear features. Furthermore, we conduct the first experimental validation of the reconstruction methodology by accurately predicting independent, downstream experimental measurements of the beam phase space, such as a longitudinal phase space (LPS) and transverse-energy correlations at the end of the beamline (located approximately 27 m from the reconstruction region) excluded from the training process. These downstream observations insofar are the best evidence of GPSR predictive accuracy, revealing that the reconstructed phase space closely approximates the ground truth physical distribution. This work shows that the reconstruction framework enables efficient, high-fidelity 6-dimensional phase space diagnostics using only standard elements, and establishes a foundation for predictive, uncertainty-aware diagnostics in a wide range of accelerator facilities.
@@ I.1 GPSR methodology with standard beamline elements
The key insight of the methodology can be conceptualized as a decomposition of the full reconstruction process into several components, each informed by physical measurements along the beamline. First, the 4-dimensional transverse phase space $(x,x^{\prime},y,y^{\prime})_{i}$ is reconstructed using quadrupole scans and transverse beam distributions measured at a screen along the straight beamline [40] as shown in Fig. 1c. Here, we note that 6-dimensional phase space of the beam is described by $(x,x^{\prime},y,y^{\prime},z,\delta)$ with $x^{\prime}=p_{x}/p_{z}$, $y^{\prime}=p_{y}/p_{z}$ where $p_{x,y,z}$ indicates momentum in $(x,y,z)$ plane, and $\delta=\left(p_{z}-p_{z,ref}\right)/p_{z,ref}$, respectively. Subscripts $i$ and $f$ denote the initial and final states. The reconstructed transverse phase space is then treated as known input for dispersive measurements. Accordingly, the horizontal beam distribution $x_{f}$ observed at the same screen with the spectrometer activated allows inference of the fractional energy spread $\delta_{i}$ through the dispersive measurement, as described in Eq. (1).
@@
In particular, since the 5-dimensional phase space is explicitly determined by measurements, the reconstruction of the longitudinal coordinate $z_{i}$ is strongly constrained by the intrinsic $(z,\delta)$ correlations described in Eq. (2). Thus, this reconstruction process enables recovery of additional correlations associated with $z_{i}$, yielding a nearly unique solution that closely approximates the 6-dimensional phase space of the beam. Thus, it has direct implications regarding prediction accuracy and uncertainty quantification. We will investigate these factors with independent downstream measurements in the following sections.
@@ I.2 Experimental setup
A schematic of the beamline elements used in the demonstration is shown in Fig. 1c. We used an X-band RF accelerating cavity (XLIN) to obtain controlled energy modulations [51], which is typically used to linearize the non-linear energy chirp [52] in XFEL facilities. For the transverse phase space rotation, we used Q1 quadrupole magnet. We used the first two dipole magnets in the bunch compressor (BC) section as the spectrometer. Particularly, the BC is mounted on the movable stage support, allowing it to be configured into a straight line [53]. In this configuration, we can measure the projected $(x,y)$ beam distributions at the straight and dispersive sections with the YAG1 screen as shown in Fig. 1c-d (See Methods for more details on the experimental setup).
Downstream of the GPSR section comprises 3 quadrupole magnets, S-band TCAV, spectrometer magnet, and the YAG2 screen. This section is mainly used to measure the longitudinal phase space and other position-momentum correlations to investigate the prediction accuracy with independent downstream measurements. It should be highlighted that that all downstream measurements are not included in the training process and used solely for validation. Total length of the beamline from the Q1 to YAG2 is roughly 27 m.
Figure 2: Comparison between measured beam distributions and predictions at the YAG1 screen. Training dataset is highlighted with orange box. White(black) contour line indicates 90th(50th) percentile, where solid(dashed) lines represent the measurement(prediction). Background image shows the measurement for a, while it illustrates the prediction for b.
In the data acquisition, we sampled beam distributions using a combination of 16 Q1 magnet strengths and 8 XLIN phase settings (see Methods for more details of the setup). An additional setting with the XLIN cavity turned off was included to obtain the initial $(z,\delta)$ correlations of the beam before manipulation by the XLIN cavity. Accordingly, there are a total of 9 XLIN phases. For each beamline configuration, we took 15 shots at a repetition rate of 2 Hz with a charge window within ±$\pm 3$% of the nominal bunch charge of 250 pC. The measured images were then averaged, and then resulting beam distribution was used as a data sample for the reconstruction process. The total number of data samples was $16\times 9\times 2=288$, where the factor of 2 corresponds to the spectrometer configurations (BC ON/OFF). It was then split in half to create a training dataset (144 samples) which was used to reconstruct the beam distribution, and a test dataset used to validate the accuracy of reconstruction predictions.
@@
Initial training was performed using the full set of 144 training samples with 2,400 iterations (See Methods for details of the reconstruction setup). We subsequently confirmed that the comparable GPSR performance can also be achieved using a reduced dataset and training iterations; it took approximately 30 minutes for 1,000 iterations on an NVIDIA A100 GPU at NERSC [56], where we used the dataset from only 5 XLIN phases (80 samples).
Figure 3: Comparison of LPSs and $(y,\delta)$ correlations at the YAG2 screen. In case of the LPS, the beam head is placed at positive $z$ while the tail is at negative $z$. Background image indicates the measured distribution. The average kinetic energy is indicated in each sub-figure, as the $\delta$ is distributed with respect to zero. White dashed(yellow solid), projected histogram indicates predicted(measured) density distribution. White(black) contour line indicates 90th(50th) percentile. Top row a-e: LPS where BC ON. Middle row f-j: LPS where BC OFF. Bottom row k-o: $(y,\delta)$ correlations with BC OFF.
@@ I.3 Prediction of training and test datasets from the reconstructed phase space
First, Fig. 2 shows a comparison between experimental measurements and reconstruction predictions for a subset of training and test settings at the YAG1 screen. In this figure, the background image in a(b) represents the measurement(prediction) [see Supplementary Figs. 1-2 for comparisons using the full datasets]. In case of the dispersive measurements with the XLIN turned off (top row), small discrepancies are observed between the predicted and measured distributions, likely due to the low signal-to-noise ratios (SNR) or RF phase drift of accelerating cavities upstream of the demonstration section. However, these small deviations do not compromise the ability of the trained model to capture detailed features of the beam, and they further demonstrate the robustness of the methodology in the presence of measurement noise.
@@ I.4 Validation of the methodology via independent downstream measurements
In addition to analyzing the training and test datasets, a critical component of validating the methodology is the comparison of reconstructed beam characteristics with independent downstream measurements that are completely excluded from the training process. We propagated the reconstructed 6-dimensional distribution through the remainder of the beamline using the elegant simulation code, which takes into account collective and nonlinear optical effects such as coherent synchrotron radiation and chromatic aberrations. The resulting predicted beam distributions were then compared with experimental measurements at the YAG2 screen, located 27 m downstream from the reconstruction point. These measurements were obtained using a combination of quadrupole magnets, S-band TCAV, and spectrometer dipole as shown in Fig. 1c.
Figure 3 presents a comparison between simulated predictions and experimental measurements of the longitudinal phase space and $(y,\delta)$ correlations under a variety of beamline conditions (see Methods for details on YAG2 measurements and Supplementary Fig. 3 for additional comparisons). The image observed at the YAG2 screen represents the projected $(z,\delta)$ distribution in $(x,y)$ plane by means of the TCAV and spectrometer magnet. This implies that focusing effects by quadrupole magnets located around the diagnostic region introduce additional transverse correlations on the beam, distorting the projected $(z,\delta)$ mapping. Thus, the measured image is a result of both longitudinal and transverse phase space projections.
@@
Nevertheless, these results confirm that the reconstruction methodology is able to accurately predict independent downstream measurements, demonstrating not only the robustness of the predictive diagnostics, but also that the initially reconstructed distribution can be considered close approximation to the actual physical beam. Thus, this validation indicates that predictive diagnostics are applicable at any beamline locations along the reconstruction section. In addition, predictive diagnostics can be extended to further downstream positions with the calibrated simulation lattice which is almost identical to the experimental condition.
Figure 4: a. Reconstructed 6-dimensional phase space. Blue histogram indicates the 1-dimensional histogram in each coordinate. b. 3-dimensional density plots of reconstructed beam phase spaces.

### v3

@@ Abstract
Particle accelerator-based user facilities are indispensable tools that power scientific discovery in the biological, chemical and material sciences [1, 2]. Understanding the precise distribution of particles in an accelerated beam is therefore critical to improving and maintaining optimal performance of these facilities. For example, X-ray Free Electron Laser (XFEL) facilities are strongly dependent on the distribution of electrons in 6-dimensional position-momentum space to generate high-brightness, ultra-short X-ray pulses [3, 4, 5, 6, 7]. The precise structure of the beam distribution during the acceleration and compression stages of the accelerator impacts overall beam quality due to collective effects such as space charge and coherent synchrotron radiation (CSR) [8, 9, 10, 11], that in-turn, negatively impacts FEL performance. As a result, methods are needed to efficiently and precisely characterize 6-dimensional phase space distribution at multiple locations along the beamline.
@@
While methods for measuring lower dimensional phase space distribution structure of particle beams are well understood, either through direct measurements including sub-sampling portions of the beam distribution [12] or via indirect measurements such as tomographic manipulations [13, 14, 15], their applicability becomes limited on 5- [16, 17] or 6-dimensional [18] phase space distributions due to largely increased number of samples required [19]. Previous attempts at characterizing the high-dimensional phase space distribution of beams require a substantial amount of dedicated experimental beam time to conduct (>>24 hrs for 5- and 6-dimensional characterization [17, 18]), making these methods impractical for use at user facilities. Additionally, these methods also require the use of specialized diagnostic elements to characterize the beam distribution, such as movable slits or transverse deflecting cavities (TCAVs) [20, 21]. These additional requirements can add complexity and operational costs to diagnose the beam at multiple locations and reduce high-value space in the accelerator beamline.
Recently, a new machine learning technique, known as generative phase space reconstruction (GPSR), has been developed to address the limitations of conventional beam diagnostic methods [22]. This approach is different from virtual diagnostics which have also been actively investigated [23, 24, 25, 26, 27, 28, 29]. The main difference between the GPSR method and virtual diagnostics is that it reconstructs the input beam distribution rather than predicting the output for a new set of input parameters. For the case of the generative model for phase space diagnostics, studies on normalizing flows [30, 31] are also actively carried out in recent years.
@@
A schematic of the GPSR technique is as follows [see Fig. 1a]: the initial 6-dimensional phase space distribution is parameterized by a generative machine learning model that is able to flexibly generate a wide variety of distribution structures. The initially generated beam is then propagated through a simulated version of the beamline using a differentiable particle tracking software [32]. To reconstruct the beam distribution, the parameters of the generative model are iteratively updated via gradient descent optimization [33] to minimize the difference between experimental measurements of the $(x,y)$ beam distribution on intercepting screens and corresponding simulated predictions. Minimizing the difference ultimately yields a reconstructed beam phase space distribution that reproduces experimental measurements.
Figure 1: a. Schematics of the GPSR method [22]. b: Cartoon showing an evolution of beam energy chirp by the accelerating field. c. Schematic view of the first bunch compressor [34] section at the PAL-XFEL. Training and test datasets are measured at the screen 1 [35, 36]. The lattice used for the reconstruction training process is in the dashed orange box. d. Validation section. The longitudinal phase space is measured at screen 2 using S-band TCAV and spectrometer [37] [e. measured longitudinal phase space for BC and XLIN OFF]. Measured beam charge and reference energy for the reconstruction process are 250 pC and 260.5 MeV, respectively.
@@
Previous experimental demonstrations of the GPSR method [22, 38, 39] have shown accurate reconstruction of the phase space distribution validated by prediction of training and test datasets and comparison of beam parameters measured by conventional diagnostics. Nonetheless, its broader applications at the accelerator facilities remains limited, as it still relies on specialized diagnostic elements such as TCAVs to rotate and probe the beam distribution in 6-dimensional phase space. In addition, prior studies have primarily focused on validating reconstructions at a specific diagnostic point, leaving predictive accuracy and uncertainty of the reconstruction at additional downstream locations unexplored.
@@
In this work, we demonstrate the first implementation and validation of the GPSR methodology for complete 6-dimensional phase space reconstruction using standard beamline elements commonly available at accelerator facilities [2, 40, 41], including accelerating cavities and spectrometer magnets. Through both simulation and experiment at the user facility, Pohang Accelerator Laboratory X-ray Free Electron Laser (PAL-XFEL) [42, 6, 7], we show that a combination of quadrupole scans, dispersive measurements, and the accelerating cavity phase scans provides sufficient information to reconstruct the detailed 6-dimensional phase space distribution of the beam including complex and nonlinear features. Furthermore, we conduct the first experimental validation of the reconstruction methodology by accurately predicting independent, downstream experimental measurements of the beam phase space, such as a longitudinal phase space (LPS) and transverse-energy correlations at the end of the beamline (located approximately 27 m from the reconstruction region) excluded from the training process. These downstream observations insofar are the best evidence of GPSR predictive accuracy, revealing that the reconstructed phase space closely approximates the ground truth physical distribution. This work shows that the reconstruction framework enables efficient, high-fidelity 6-dimensional phase space diagnostics using only standard elements, and establishes a foundation for predictive diagnostics in a wide range of accelerator facilities.
@@ I.1 GPSR methodology with standard beamline elements
The key insight of the methodology can be conceptualized as a decomposition of the full reconstruction process into several components, each informed by physical measurements along the beamline. First, the 4-dimensional transverse phase space $(x,x^{\prime},y,y^{\prime})_{i}$ is reconstructed using quadrupole scans and transverse beam distributions measured at a screen along the straight beamline [22]. Here, we note that 6-dimensional phase space of the beam is described by $(x,x^{\prime},y,y^{\prime},z,\delta)$ with $x^{\prime}=p_{x}/p_{z}$, $y^{\prime}=p_{y}/p_{z}$ where $p_{x,y,z}$ indicates normalized momentum in $(x,y,z)$ plane; $\gamma m_{e}\beta_{x,y,z}c$ normalized by $m_{e}c^{2}$ where $m_{e}$ indicates the electron mass. $\delta=\left(p_{z}-p_{z,ref}\right)/p_{z,ref}$ where $p_{z,ref}$ is the reference particle’s longitudinal momentum. Subscripts $i$ and $f$ denote initial and final states, associated with the beam before and after the propagation to the beamline. The reconstructed transverse phase space is then predominated and treated as constraints for dispersive measurements. Accordingly, the horizontal beam distribution $x_{f}$ observed at the same screen with the spectrometer activated allows inference of the fractional energy spread $\delta_{i}$ through the dispersive measurement, as described in Eq. (1).
@@
In particular, since the 5-dimensional phase space is explicitly determined by measurements, the reconstruction of the longitudinal coordinate $z_{i}$ is strongly constrained by the intrinsic $(z,\delta)$ correlations described in Eq. (2). Thus, this reconstruction process enables recovery of additional correlations associated with $z_{i}$, yielding a nearly unique solution that closely approximates the 6-dimensional phase space of the beam. Thus, it has direct implications regarding prediction accuracy and uncertainty of the reconstruction. We will investigate these factors with independent downstream measurements in the following sections.
@@ I.2 Experimental setup
A schematic of the beamline elements used in the demonstration is shown in Fig. 1c. We used an X-band RF accelerating cavity (XLIN) to obtain controlled energy modulations [46], which is typically used to linearize the non-linear energy chirp [47] in XFEL facilities. For the transverse phase space rotation, we used Q1 quadrupole magnet. We used the first two dipole magnets in the bunch compressor (BC) section as the spectrometer. Particularly, the BC is mounted on the movable stage support, allowing it to be configured into a straight line [48]. In this configuration, we can measure the projected $(x,y)$ beam distributions at the straight and dispersive sections with the screen 1 (See Methods for more details on the experimental setup).
Downstream of the GPSR section comprises 3 quadrupole magnets, S-band TCAV, spectrometer magnet, and the screen. This section is mainly used to measure the longitudinal phase space and other position-momentum correlations to investigate the prediction accuracy with independent downstream measurements. It should be highlighted that that all downstream measurements are not included in the training process and used solely for validation. Total length of the beamline from the Q1 to screen 2 is roughly 27 m.
Figure 2: Comparison between measured beam distributions and predictions at the screen 1. Training dataset is highlighted with orange box. $k$ indicates the quadrupole magnetic field strength in T/m. Contour lines indicate 90 and 50th percentiles, where solid(dashed) lines represent the measurement(prediction). Background image shows the measurement for a, while it illustrates the prediction for b.
In the data acquisition, we sampled beam distributions using a combination of 16 Q1 magnet strengths and 8 XLIN phase settings (see Methods for more details of the setup). An additional setting with the XLIN cavity turned off was included to obtain the initial $(z,\delta)$ correlations of the beam before manipulation by the XLIN cavity. Accordingly, there are a total of 9 XLIN phases. For each beamline configuration, we took 15 shots at a repetition rate of 2 Hz with a charge window within $\pm 3$% of the nominal bunch charge of 250 pC. The measured images were then averaged, and then resulting beam distribution was used as a data sample for the reconstruction process. The total number of data samples was $16\times 9\times 2=288$, where the factor of 2 corresponds to the spectrometer configurations (BC ON/OFF). It was then split in half to create a training dataset (144 samples) which was used to reconstruct the beam distribution, and a test dataset used to validate the accuracy of reconstruction predictions.
@@
Initial training was performed using the full set of 144 training samples with 2,400 iterations (See Methods for details of the reconstruction setup). We subsequently confirmed that the comparable GPSR performance can also be achieved using a reduced dataset and training iterations; it took approximately 30 minutes for 1,000 iterations on an NVIDIA A100 GPU at NERSC [51], where we used the dataset from only 5 XLIN phases (80 samples). Accordingly, we will show the reconstruction results using reduced number of data samples.
Figure 3: Comparison of LPSs and $(y,\delta)$ correlations at the screen 2. In case of the LPS, the beam head is placed at positive $z$ while the tail is at negative $z$. Background image indicates the measured distribution. The average kinetic energy is indicated in each sub-figure, as the $\delta$ is distributed with respect to zero. White dashed(yellow solid), projected histogram indicates predicted(measured) density distribution. White contour lines indicate 90 and 50th percentiles. Top row a-e: LPS where BC ON. Middle row f-j: LPS where BC OFF. Bottom row k-o: $(y,\delta)$ correlations with BC OFF.
@@ I.3 Prediction of training and test datasets from the reconstructed phase space
First, Fig. 2 shows a comparison between experimental measurements and reconstruction predictions for a subset of training and test settings at the screen 1. In this figure, the background image in a(b) represents the measurement(prediction) [see Supplementary Figs. 1-2 for comparisons of full datasets]. In case of the dispersive measurements with the XLIN turned off (top row), small discrepancies are observed between the predicted and measured distributions, likely due to the low signal-to-noise ratios (SNR) or RF phase drift of accelerating cavities upstream of the demonstration section. However, these small deviations do not compromise the ability of the trained model to capture detailed features of the beam, and they further demonstrate the robustness of the methodology in the presence of measurement noise.
@@ I.4 Validation of the methodology via independent downstream measurements
In addition to analyzing the training and test datasets, a critical component of validating the methodology is the comparison of reconstructed beam characteristics with independent downstream measurements that are completely excluded from the training process. We propagated the reconstructed 6-dimensional distribution through the remainder of the beamline using the elegant simulation code, which takes into account collective and nonlinear optical effects such as coherent synchrotron radiation and chromatic aberrations. The resulting predicted beam distributions were then compared with experimental measurements at the screen 2, located 27 m downstream from the reconstruction point. These measurements were obtained using a combination of quadrupole magnets, S-band TCAV, and spectrometer dipole as shown in Fig. 1c.
Figure 3 presents a comparison between simulated predictions and experimental measurements of the longitudinal phase space and $(y,\delta)$ correlations under a variety of beamline conditions (see Methods for details on screen 2 measurements and Supplementary Fig. 3 for additional comparisons). The image observed at the screen 2 represents the projected $(z,\delta)$ distribution in $(x,y)$ plane by means of the TCAV and spectrometer magnet. This implies that focusing effects by quadrupole magnets located around the diagnostic region introduce additional transverse correlations on the beam, distorting the projected $(z,\delta)$ mapping. Thus, the measured image is a result of both longitudinal and transverse phase space projections.
@@
Nevertheless, these results confirm that the reconstruction methodology is able to accurately predict independent downstream measurements, demonstrating not only the robustness of the predictive diagnostics, but also that the initially reconstructed distribution can be considered close approximation to the actual physical beam. Thus, this validation indicates that predictive diagnostics are applicable at any beamline locations along the reconstruction section. In addition, the reconstructed phase space information can be used either as a target or as an input to calibrate the simulated model of the upstream beamline and further downstream lattice for the predictive diagnostics.
Figure 4: a. Reconstructed 6-dimensional phase space. Contour lines indicate 90 and 50th percentile, while different line styles represent three cases in the error study. ”Reference” indicates the nominal case. Note that only 90th percentile is shown for $(z-\delta)$ distribution. Blue histogram indicates the 1-dimensional histogram in each coordinate. b. 3-dimensional scatter/density plot of the reconstructed beam phase space.